OpenStudy (anonymous):

Nathan has 80 stamps in his collection. He adds 1 stamp to it today. Each day he plans to add twice the number of stamps as the previous day. If he keeps adding stamps at this rate for n days, which recursive function represents the number of stamps he has on any day in the future?

OpenStudy (anonymous):

The recursive function is: A) next=now+2^n B)next=now+2^n-1 C)next=now-2^n Starting at A) 80 B)2^n C)80^n